Determining if allele frequencies are changing from one generation to the next (microevolution) from the number of individuals of each genotype present:

The following steps are used to determine if allele frequencies are changing:

  1. Calculate Allele frequency from the number of individuals of each genotype

  2. Calculate expected genotypic frequencies and individuals in a population from allele frequencies:

  3. Test the goodness of fit between the data and the Hardy Weinberg equilibrium model generated expectations.

The following problems are the calculations used to determine if allele frequencies are changing-if the gene is not in equilibrium. The first two sets will have you calculate the steps used to detect allele changes. The final set of problems will use all of the steps to determine if allele frequencies are changing in a population.

  1. Calculating Allele frequency from the number of individuals of each genotype

2.1 In a species of lizard, individuals have either thirteen or fourteen brown and white horizontal stripes on their back. This trait is controlled by genes at a single locus, with the allele for thirteen stripes dominant to the allele for fourteen. A population of lizards is sampled; 91 with thirteen stripes and 37 with fourteen stripes are recorded. Of the 91 with thirteen stripes, 29 are homozygous dominant. What are the frequencies of the dominant and recessive alleles observed in this sample?

p =

q =

2.2 In a species of bird, individuals have either red or yellow feathers. This trait is controlled by genes at a single locus, with the red allele dominant to the yellow allele. A population of birds is sampled; 15 red and 9 yellow are recorded. Of the 15 red ones, 7 are homozygous dominant. What are the frequencies of the dominant and recessive alleles observed in this sample?

p =

q =

2.3 In humans, individuals have either a dimpled chin or a chin without a dimple. This trait is controlled by genes at a single locus, with the allele for a dimple dominant to the allele to not have a dimple. A population of humans is sampled; 24 have dimples and 14 do not. 11 of the people with dimples are homozygous dominant. What are the frequencies of the dominant and recessive alleles observed in this sample?

p = _____

q = _____

2.4 In a species of pine tree, individuals have either narrow or wide cones. This trait is controlled by genes at a single locus, with the narrow cone allele dominant to the wide cone allele. A forest is sampled; 112 trees with narrow cones and 126 trees with wide cones are counted. Of the narrow cone trees, 55 are homozygous dominant. What are the frequencies of the dominant and recessive alleles observed in this sample?

p = _____

q = _____

2.5 In humans, individuals may or may not have a widow’s peak. This trait is controlled by genes at a single locus, with the allele for having a widow’s peak dominant to the allele for not having one. A population of humans is sampled; 41 have a widow’s peak and 59 do not. 12 of the people with a widow’s peak are homozygous dominant. What are the frequencies of the dominant and recessive alleles observed in this sample?

p = _____

q = _____


2.6 In a species of corn, individuals have either long or short corn cobs. This trait is controlled by genes at a single locus, with the long cob allele dominant to the short cob allele. A farmer counts 105 long cob and 149 short cob plants in a field. Of the long cob plants, 81 are homozygous dominant. What are the frequencies of the dominant and recessive alleles observed in this sample?

p = _____

q = _____
Answer #1

Hi, Here I have solved all the p and q values for every given problem, I hope you can solve other statements based on these values, in case of any problem, you may please mention in the comment box .

Thanks

solution - Consideren 13 Stecip = 513 = & À for the 14 Strip = S14 = (a Solving Homozygous dominant 13 Strips AA) S13 S13 = 2Allelic fuequencies) P= f(A) = f (AA) + f (49) = 0.226 +5x0.484 > 0, 226 + 0,242 50.468 *9=f(q) = f (99) +%sf (49 3 0 289 +19Consider Red = R Yellow a r Homozygous dominant (RR) = 7 Heterozygous total .. (Rr) Homozygous viccessive! (rr) = 9 Gremo typ1 2 Allelic Fuequencies p=f(Rfg = f(RR) + (RA) - = 0.28 +636) -0.28 +0.18 20.46 q=f(21) = 0.36 +12(0.36) = 0.36 +0.18 - 0.54og (Genotypic fereq qencuest FCDD) - 110428,5 f (Dd) = 13 = 0.342 f (dd) = 14 = 0. 368 Auelie frequencies pe f (DJ 20.289 + 16.4 total Consider a Narrow cone A. wide cone = a Норлогодио ( АЯ) = 55 1 Heterozygou (Aq) = 57 Homozygous (99) = 426 J GenotLoo 100 Genotypic feequencies f (PP) = 6-12 =6.12 f (PP) = 29-(0.29 OCO fcpp) = 59.0.593 Allelic Frequencies p= f (P) = 0.1278.6 Consider Long Cob = Short Cobal Номозоча {LL ) = 81 неtvло ( L4) = 24 254 Келмо газе» (CL) з 149 ) Genotypic frequenejesAllelio frequencies P= f(t) = 0.318+(0-094) = 0.318+ 0.047 ĉ 0.365 9. f 4) = 0.586 + yg (0.094) © 0.586 +0.047 = 0.633 Pa 0.3